OpenStudy (anonymous):

What is the solution to the proportion 6/10 = 9/x A. 30 B. 13 C. 15 D. 5.4

OpenStudy (anonymous):

First Let's solve your equation step-by-step. 6/10=9/x Step 1: Cross-multiply. 6/10=9/x 6*x=(9)*(10) 6x=90 Step 2: Divide both sides by 6. 6x/6=90/6 x=15 Answer: x=15

OpenStudy (anonymous):

Did that help

OpenStudy (anonymous):

Yes very much

OpenStudy (anonymous):

I think i got the others now

OpenStudy (anonymous):

Ill tag you again if i think my answer needs to be checked on another question

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Just tag and I will be more than glad to help